# CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners

### v1

@@ 3.1 Rethinking KE from the Circuit View
Despite the success of current KE on single facts benchmarks, from previous studies Zhang et al. (2024d); Zhong et al. (2023), and our analysis in §4, we can see that the edited model’s performance on multi-hop reasoning tasks is often unsatisfactory. Building on our previously identified circuit for multi-hop reasoning, we rethink the reason why current knowledge editing methods fail under multi-hop reasoning circumstances.
@@
When updating a piece of knowledge $(e,r,o\rightarrow o^{\prime})$, the most popular knowledge editing techniques would modify the parameters that are responsible for the knowledge. There are two kinds of paradigms: editing the Feed-Forward Networks (FFN) in the early layers, such as ROME Meng et al. (2022) and MEMIT Meng et al. (2023) or modifying the later layers’ FFN output, like WISE Wang et al. (2024c) and T-Patcher Huang et al. (2023). This is mainly based on the key-value memory features of the FFN Geva et al. (2020). However, some studies have queried the effectiveness of these localization settings Chang et al. (2024); Hase et al. (2024) as the localization area is not correlated to the performance of the knowledge editing methods. Here, we propose a unified view of the mechanisms and limitations from the circuit perspective. From Figure 2 (c), these two editing paradigms can be achieved by a gated function $\mathcal{G(\mathbf{x})}$:
@@
When the gating function $\mathcal{G}(\mathbf{x})$ is activated by the input $\mathbf{x}$, the edit term $\bm{\delta}(\mathbf{x})$ is applied, thereby modifying the knowledge within the computational circuit. ROME-style would modify the weight $\bm{W}$ with a perturbation $\bm{\Delta}$ and obtain a new weight $\bm{W}^{\prime}=\bm{W}+\bm{\Delta}$. Here, $\bm{\delta}(\mathbf{x})=\bm{\Delta}\mathbf{x}$. When calculating the $\Delta$, ROME-style methods, apply the least squares estimation and null space constraint to make sure the $\Delta$ is only activated by the corresponding entity representation $\bm{e}_{in}$ and keep the original output for other representations. In parallel, WISE-style editing methods would directly introduce the new weight $\bm{W}^{\prime}$ that would be activated by the related representation $\bm{e}_{in}$, and $\bm{W}^{\prime}$ would encode the updated knowledge. Here, $\bm{\delta}(\mathbf{x})=(\bm{W}^{\prime}-\bm{W})\mathbf{x}$.
@@
To see the editing mechanism better, we first compare the rank of the target answer at the last token position via MEMIT, WISE, and the original model in Figure 4. We edit the fact: ‘The official language of Japan is Japanese $\rightarrow$ Korean.’ From the figure, we can see that in the original model and MEMIT-edited model, the answer token is dealt with gradually through the mid-to-later layers, and MEMIT would make this happen in advance. The computation of the ROME-style method for the new knowledge recall is: $F(\tilde{e},r)=o^{\prime},\tilde{e}=e+\bm{\Delta}e$, which modifies the knowledge stored in the previous layer and gives us the new representation $e^{\prime}$ for further computing. On the contrary, the WISE method would directly alter the information at the edited layer as we can see the sharp drop at layer 29. In particular, the editing would take effect when the added or updated parameters are activated by query representation and work as $F(e,r)=o^{\prime}$.
@@
H-Acc.$\uparrow$ MAcc.$\uparrow$ H-Acc.$\uparrow$ MAcc.$\uparrow$ H-Acc.$\uparrow$ MAcc.$\uparrow$
Pre-edited LLaMA3-8B-Ins 79.0 27.0 78.4 28.6 71.0 5.3
LoRA 66.0 27.6 64.7 24.6 92.3 66.0
@@
In single-hop knowledge editing, these kinds of methods would give us the correct information, but for the multi-hop cases, this would fail. As shown in Figure 2 (b), both these layer-specific editing methods cannot propagate the updated knowledge to the reasoning circuit, leading to unsatisfactory multi-hop reasoning performances. An essential requirement for these methods is that the gated function $\mathcal{G(\mathbf{x})}$ is activated by the specific representation $\bm{e}_{in}$. However, under the multi-hop reasoning scenario, the model would deal with different single-hop questions in different layers, like the two-hop reasoning circuit in §2: for ROME-style editing, if the new fact $(e,r,o\rightarrow o^{\prime})$ is the second-hop question and the entity $e$ appears after the edited layers, the gated function would $\mathcal{G(\mathbf{x})}$ not be activated and the model would still follow the previous stale knowledge $F(e,r)$ instead of $F(\tilde{e},r)$ and give us the wrong answer. Likewise, the WISE-style editing would retain reliance on the original knowledge when the new fact is finished in the former layers as the first hop, bypassing the edit function in later layers and cascading the error in the subsequent reasoning. Moreover, we conduct experiments to view the limitations better, and Table 8 show that editing the first-hop in WISE leads to poorer performance compared to editing the second-hop. Conversely, when using ROME to edit the second-hop, the performance is worse than when editing the first-hop. In conclusion, these layer-specific editing methods cannot learn the new knowledge generally to make the knowledge usable in downstream reasoning tasks.
3.2 Proposed Method: CaKE
Inspired by previous analysis, we propose a novel method, Circuit-aware Knowledge Editing (CaKE), which enhances the model’s ability to update and effectively utilize knowledge. CaKE comprises two key components: (1) generating circuit-aware training data that explicitly requires reasoning with the updated knowledge, and (2) training the model to construct robust reasoning circuits that integrate the new knowledge.
Data Generation
To ensure that the model builds effective reasoning circuits, we address two critical challenges: preventing failure propagation and mitigating weak signals (as identified in §2), while ensuring that updated knowledge is properly integrated across different layers (as described in §3.1).
Method Model MQUAKE-CF MQUAKE-CF-v2 MQUAKE-T
H-Acc.$\uparrow$ MAcc.$\uparrow$ H-Acc.$\uparrow$ MAcc.$\uparrow$ Hop-wise.$\uparrow$ MAcc.$\uparrow$
Pre-edited Qwen2.5-7B-Ins 73.4 40.7 72.8 39.5 56.1 15.6
LoRA 35.1 24.9 36.5 25.9 25.0 28.6
@@ 3.2 Proposed Method: CaKE
[…]
For each updated knowledge item, we construct the following contexts to mitigate these issues: (1) Original Narrative: We begin by generating straightforward factual statements that explicitly convey the updated information. For example, when updating the fact $k$: (PersonX, citizen_country, Switzerland → Japan), we use the narrative representation: ‘PersonX is a citizen of Japan’ and generate several paraphrases. These statements serve as the foundation for the model to learn the updated knowledge. (2) Circuit-aware Tasks: Next, we design specialized reasoning scenarios that address the identified circuit-level challenges, as illustrated in Figure 2(c). Moreover, to avoid introducing extraneous knowledge that could leak into downstream evaluations—and to test the generalization of our method (inspired by prior research Zhang et al. (2024c))—we incorporate ad-hoc features into these scenarios. These tasks link the facts with intermediate attributes or reasoning steps and fall into two categories: Late-layer Knowledge Integration: These tasks ensure that the updated knowledge is effectively learned in the later layers, alleviating issues such as weak signals and the limitations of ROME-style editing. For the fact $k$, we construct prompts like: ‘Suppose {random_entity_1} wears red clothes, {random_entity_2} wears blue clothes, and {PersonX} wears green clothes. The country of citizenship of the person in green is:’ Here, the model is expected to output ‘Japan,’ requiring it to employ the new fact $k$ in later layers. Reasoning Circuit Enhancement: These tasks require the model to use the updated knowledge for subsequent reasoning, thereby mitigating propagation failure, weak signal, and WISE-style’s limitations. Following the same fact $k$: ‘In a book about countries, Japan is mentioned on page 6 of the book, while China is mentioned on page 72. On which page of the book is the country of citizenship of the {PersonX} shown?’ Here, the model must first recall the updated citizenship (Japan) and then use this information to determine the correct page number (6). Furthermore, for each knowledge type, we develop specific task templates and leverage GLM-4-plus GLM et al. (2024) to generate data using randomly selected related entities (detailed in Appendix A).
Edit Training
After obtaining the curated circuit-aware data $\mathcal{D}$, we fine-tune the LLM using LoRA across all layers, enabling the model to optimize its internal knowledge organization. We minimize the cross-entropy loss $\mathcal{L}$ between the model’s outputs and the ground-truth tokens expressing the updated fact:
4 Experiments
@@ 4.1 Experiment Settings
We consider several knowledge editing baselines, including: IFMET Zhang et al. (2024d), AlphaEdit Fang et al. (2024), ROME Meng et al. (2022), MEMIT Meng et al. (2023),WISE Wang et al. (2024c) and MeLLo Zhong et al. (2023). Here, AlphaEdit, ROME, and MEMIT are methods that edit the model’s parameters at early layers; WISE adds additional parameters at later layers, and IFMET edits both the early and later layers’s FFN to achieve better multi-hop reasoning performance. MeLLo is a prompt-based retrieval-augmented method. We conduct experiments on LLAMA-3-8B-Instruct, Qwen-2.5-7B-Instruct, and LLAMA-3-70B-Instruct.
@@
Following Zhong et al. (2023), we evaluate model performance using Multi-hop Accuracy (MAcc) and Hop-wise Answering Accuracy (H-Acc). MAcc measures the accuracy of multi-hop question answering, while H-Acc assesses correctness at each reasoning step. Higher values indicate better performance. For KE, we also need to consider locality, which ensures edits do not affect unrelated knowledge and abilities. To assess this, we evaluate the model on general benchmarks, including CommonsenseQA Talmor et al. (2019), BigBenchHard Suzgun et al. (2023), MMLU Hendrycks et al. (2021), and GSM8k Cobbe et al. (2021).
@@ 4.2 Experiments Results
Table 3 and Table 4 summarizes our results. Although current KE methods achieve high hop-wise accuracy (H-Acc.), their performance on the three versions of MQuAKE is quite low (with an average accuracy of less than 20%). For example, MEMIT and ROME achieve over 80% accuracy on single-hop questions in MQuAKE-v2; however, their accuracy on multi-hop reasoning drops to only around 10%, indicating that the LLM fails to effectively utilize the updated knowledge during reasoning. In contrast, CaKE demonstrates significant improvements in multi-hop reasoning. On the LLAMA3-8B-Instruct model, CaKE achieves accuracies of 57.3, 57.2, and 81.5 on MQuAKE-CF, MQuAKE-CF-v2, and MQuAKE-T, respectively—outperforming all compared methods. Additionally, IFMET, which also considers different layers for multi-hop reasoning but neglects the information flow within the circuit, performs not as well as CaKE. Moreover, when compared with RAG-based methods such as MeLLo, CaKE also yields better results. Furthermore, compared to the baseline LoRA tuning methods that simply incorporate the raw knowledge, the improvements observed with CaKE underscore the effectiveness of our approach. Interestingly, while LoRA demonstrates strong performance on the 70B model, reflecting enhanced learning capabilities in larger models, CaKE still achieves an additional improvement of approximately 10%. Results in Qwen-2.5-Instruct also demonstrate the same phenomenon.
Locality Performance
@@ 5 Analysis
5.1 Position and Number of Hop
We then examine the effects of the number of edits and the position of the updated knowledge in multi-hop scenarios, with results shown in Figure 5. Notably, even when the model is trained solely on two-hop questions, CaKE yields improvements across varying numbers of editing hops. The benefits are particularly pronounced for four-hop questions, where methods like IFMET (designed only for two-hop scenarios) struggle. Besides, CaKE enhances performance regardless of the position of the edited knowledge within the multi-hop questions, demonstrating the generalizability of CaKE .
@@ 5.2 Case Analysis
Figure 6 displays the logits of $e_{2}$ and $r_{2}$ at $t_{2}$ for models edited by different methods. As shown, the bridge entity $e_{2}$ in CaKE exhibits significantly stronger logits compared to those of AlphaEdit and MEMIT, indicating that CaKE effectively constructs the reasoning circuit and propagates the necessary information to the target position. Similarly, the $r_{2}$ information is more prominent in CaKE , further demonstrating its superiority in circuit construction and information flow.
5.3 Discussion with Chain-of-Thought
@@
Instead of directly providing an answer, chain-of-thought (CoT) reasoning generates intermediate steps sequentially. As proposed by Yang et al. (2024c), CoT not only facilitates knowledge activation in large language models but also transforms them into effective in-context reasoners. The CoT process builds a chain of relevant facts within the prompt context, where each step’s output serves as an in-context memory that subsequent steps can reference. This approach reduces the risk of losing track of intermediate facts as the sequence length increases, thereby promoting more coherent multi-hop reasoning. Moreover, because a significant portion of the model’s knowledge is stored in earlier layers, CoT can better leverage these neurons by decomposing complex questions into simpler sub-questions Wang et al. (2024g); Yao et al. (2025). Consequently, the reasoning circuit required for a single-hop inference is much simpler than that for multi-hop reasoning. This observation aligns with recent findings Li et al. (2024), which demonstrate that fast thinking without CoT leads to larger gradients and greater gradient disparities across layers compared to CoT. Nonetheless, inconsistencies in the intermediate reasoning steps still occur, highlighting potential areas for improvement. We believe that further analysis is needed to address these issues, and we leave this exploration for future work.

### v2

@@ 3.1 Rethinking KE from the Circuit View
Here, we aim to figure out what happens when we edit the model with the current KE methods.
@@
When updating a piece of knowledge $(e,r,o\rightarrow o^{\prime})$, the most popular knowledge editing techniques would modify the parameters that are responsible for the knowledge. There are two kinds of paradigms: editing the Feed-Forward Networks (FFN) in the early layers, such as ROME Meng et al. (2022) and MEMIT Meng et al. (2023) or modifying the later layers’ FFN output, like WISE Wang et al. (2024c) and T-Patcher Huang et al. (2023). This is mainly based on the key-value memory features of the FFN Geva et al. (2020). However, some studies have queried the effectiveness of these localization settings Chang et al. (2024); Hase et al. (2024) as the localization area is not correlated to the performance of the knowledge editing methods. Here, we propose a unified view of the mechanisms and limitations from the circuit perspective. ROME-style would modify the weight $\bm{W}$ with a perturbation $\bm{\Delta}$ and obtain a new weight $\bm{W}^{\prime}=\bm{W}+\bm{\Delta}$. When calculating the $\Delta$, ROME-style methods, apply the least squares estimation and null space constraint to make sure the $\Delta$ is only activated by the corresponding entity representation $\bm{e}_{in}$ and keep the original output for other representations. In parallel, WISE-style editing methods would directly introduce the new weight $\bm{W}^{\prime}$ that would be activated by the related representation $\bm{e}_{in}$, and $\bm{W}^{\prime}$ would encode the updated knowledge.(More details in Appendix B.2). Hence, these two editing paradigms can be represented uniformly by a gated function $\mathcal{G(\cdot)}$:
@@
[…]
@@
[…]
Pre-edited
LLaMA3-8B-Ins
79.0 27.0 78.4 28.6 71.0 5.3
AdaLoRA 66.0 27.6 64.7 24.6 92.3 66.0
@@
Table 2: Comparison of CaKE with existing methods on MQuAKE for LLAMA3-8B-Instruct and LLAMA3-70B-Instruct. Due to the computational limitations, we just ran the LoRA and MeLLo in the 70B model. The best results are highlighted in bold, while the second-best results are underlined. $\clubsuit$} means the results are based on our re-implementation since the original code is not open by the authors, and we will update it after the source code is open.
In single-hop knowledge editing, both these kinds of methods would give us the correct information, but for the multi-hop cases, they would fail. As shown in Figure 2 (b), both these layer-specific editing methods cannot propagate the updated knowledge to the reasoning circuit, leading to unsatisfactory multi-hop reasoning performances. Consider the two-hop reasoning process from §2: the model must first correctly compute $e_{2}=F_{1}(\langle e_{1},r_{1}\rangle)$ in early layers. The representation of $e_{2}$ then propagates to the final token position $t_{2}$ (typically where answers are generated), where it combines with $r_{2}$ to compute $e_{3}=F_{2}(\langle e_{2},r_{2}\rangle)$ in later layers.
WISE-style editing shows critical limitations when handling first-hop facts $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$ in multi-hop reasoning. As the edit is applied to later layers, the early layers remain unchanged and continue to produce the original $e_{2}$ representation during computation. This creates a fundamental mismatch: while the later layers perform the second-hop computation $F_{2}(\langle e_{2},r_{2}\rangle)$, they operate on the unmodified $e_{2}$ from early layers. Consequently, the gating mechanism $\mathcal{G}(\cdot)$ designed for first-hop edits becomes effectively bypassed in the reasoning process. Similarly, ROME-style editing fails when the edited fact $(e_{2},r_{2},e_{3}\rightarrow e_{3}^{\prime})$ serves as the second-hop question. For the edit to take effect, the gating function $\mathcal{G}(\cdot)$ must be activated by $e_{2}$ in early layers. However, $e_{2}$’s representation only appears after the first hop completes in the computational pathway - potentially after the edited layers. In this scenario, the gated function $\mathcal{G(\mathbf{x})}$ in earlier layers remains unactivated, causing the model to default to stale knowledge and produce incorrect answers.
3.2 Proposed Method: CaKE
Inspired by previous analysis, we propose a novel method, Circuit-aware Knowledge Editing (CaKE), which makes sure the models build the reasoning circuit with the updated knowledge. As we show in the previous section, a successful reasoning circuit is one that, after editing the model’s knowledge, ensures: The updated computation $F_{1}$ or $F_{2}$ accurately reflects the new knowledge, and the bridge entity $e_{2}$ is correctly computed and propagated to $t_{2}$. Hence, simply editing a single layer or several layers is not enough to enable the circuit for reasoning. Here, CaKE comprises two key components: (1) generating circuit-aware data that explicitly requires reasoning with the updated knowledge, and (2) training the model to construct robust reasoning circuits that integrate the new knowledge.
Data Generation
For each updated knowledge item, we construct the following contexts to mitigate these issues: (1) Original Narrative: We begin by generating straightforward factual statements that explicitly convey the updated information. For example, when updating the fact $k$: (PersonX, citizen_country, Switzerland → Japan), we use the narrative representation: ‘PersonX is a citizen of Japan’ and generate several paraphrases. These statements serve as the foundation for the model to learn the updated knowledge. (2) Circuit-aware Tasks: Next, we design specialized reasoning scenarios that address two critical challenges: preventing failure propagation and mitigating weak signals, while ensuring that updated knowledge is properly integrated across different layers (in Figure 2c). Moreover, to avoid introducing extraneous knowledge that could leak into downstream evaluations—and to test the generalization of our method (inspired by prior research Zhang et al. (2024c))—we incorporate ad-hoc features into these scenarios. Particularly, these tasks link the facts with intermediate attributes or reasoning steps and fall into two categories: Late-layer Knowledge Integration: These tasks ensure that the updated knowledge is effectively learned in the later layers, alleviating issues such as weak signals and the limitations of ROME-style editing. Take the fact $k$: (PersonX, citizen_country, Switzerland → Japan) as an example; we construct a seed prompt like: ‘Suppose {random_entity_1} wears red clothes, {random_entity_2} wears blue clothes, and {PersonX} wears green clothes. The country of citizenship of the person in green is:’ Here, the model is expected to output ‘Japan,’ requiring it to employ the new fact $k$ in later layers. Reasoning Circuit Enhancement: These tasks require the model to use the updated knowledge for subsequent reasoning, thereby mitigating propagation failure and WISE-style’s limitations. Following the same fact $k$, the seed prompt is ‘In a book about countries, Japan is mentioned on page 6 of the book, while China is mentioned on page 72. On which page of the book is the country of citizenship of the {PersonX} shown?’ Here, the model must first recall the updated citizenship (Japan) and then use this information to determine the correct page number (6).
For each relation type, we design these seed task templates and employ GLM-4-plus GLM et al. (2024) to generate diverse expressions following these templates (see Appendix A for details). Specifically, we create 3 distinct samples per category for each edited fact, which our experiments show are sufficient to enable effective reasoning with the updated knowledge. This minimal data requirement demonstrates the efficiency of our approach in adapting models to new information.
Edit Training
After obtaining the curated circuit-aware data $\mathcal{D}$, we fine-tune the LLM using LoRA, enabling the model to optimize its internal knowledge organization. We minimize the cross-entropy loss $\mathcal{L}$ between the model’s outputs and the ground-truth tokens expressing the updated fact:
4 Experiments
Method Model MQUAKE-CF MQUAKE-CF-v2 MQUAKE-T
@@ 4 Experiments
Pre-edited
Qwen2.5-7B-Ins
73.4 40.7 72.8 39.5 56.1 15.6
AdaLoRA 35.1 24.9 36.5 25.9 25.0 28.6
@@
Table 3: Comparison of CaKE with existing methods on MQuAKE on Qwen2.5-7B-Instruct. The best results are highlighted in bold, while the second-best results are underlined. $\clubsuit$} means the results are based on our own implementation since the original code is not open by the authors, and we will update it after the source code is open.
@@ 4.1 Experiment Settings
We consider several knowledge editing baselines, including: IFMET Zhang et al. (2024d), AlphaEdit Fang et al. (2024), ROME Meng et al. (2022), MEMIT Meng et al. (2023),WISE Wang et al. (2024c) and MeLLo Zhong et al. (2023). Here, AlphaEdit, ROME, and MEMIT are methods that edit the model’s parameters at early layers; WISE adds additional parameters at later layers, and IFMET edits both the early and later layers’ FFN to achieve better multi-hop reasoning performance. MeLLo is a prompt-based retrieval-augmented method that keeps the model’s parameters unchanged. We conduct experiments on LLAMA-3-8B-Instruct, Qwen-2.5-7B-Instruct, and LLAMA-3-70B-Instruct.
@@
Following Zhong et al. (2023), we evaluate model performance using Multi-hop Accuracy (MAcc) and Hop-wise Answering Accuracy (H-Acc). MAcc measures the accuracy of multi-hop question answering, while H-Acc assesses correctness at each reasoning step. For both metrics, we consider a prediction correct if the ground-truth answer appears in the generated text as Cohen et al. (2024); Zhong et al. (2023) did. Higher values indicate better reasoning capability. For KE, we also need to consider locality, which ensures edits do not affect unrelated knowledge and abilities. To assess this, we evaluate the model on general benchmarks, including CommonsenseQA Talmor et al. (2019), BigBenchHard Suzgun et al. (2023), MMLU Hendrycks et al. (2021), and GSM8k Cobbe et al. (2021).
@@ 4.2 Experiments Results
We show the results for LLAMA3-8B-Instruct in Table 2 and Qwen2.5-7B-Instruct in Table 3. From the table, we can find that although current KE methods achieve high hop-wise accuracy (H-Acc.), their performance on the three versions of MQuAKE is quite low (with an average accuracy of less than 20%). For example, MEMIT and ROME achieve over 80% accuracy on single-hop questions in MQuAKE-v2; however, their accuracy on multi-hop reasoning drops to only around 10%, indicating that the LLM fails to effectively utilize the updated knowledge during reasoning. In contrast, CaKE demonstrates significant improvements in multi-hop reasoning. In the LLAMA3-8B-Instruct model, CaKE achieves accuracies of 57.3, 57.2 and 81.5 in MQuAKE-CF, MQuAKE-CF-v2 and MQuAKE-T, respectively, outperforming all the compared methods. Additionally, IFMET, which also considers different layers for multi-hop reasoning but neglects the information flow within the circuit, performs not as well as CaKE. Moreover, when compared with RAG-based methods such as MeLLo, CaKE also yields better results. Furthermore, compared to the baseline LoRA tuning methods that simply incorporate the raw knowledge, the improvements observed with CaKE underscore the effectiveness of our approach. Results in Qwen-2.5-Instruct also demonstrate the same phenomenon.
Position and Number of Hop
We also compare the performance on different hops and positions in Figure 4. Even when the model is trained solely on two-hop questions, CaKE yields improvements across varying numbers of editing hops. The benefits are particularly pronounced for four-hop questions, where methods like IFMET (designed only for two-hop scenarios) struggle. Besides, CaKE enhances performance regardless of the position of the edited knowledge within the multi-hop questions, demonstrating the generalizability of CaKE.
Efficiency and Scalability
We evaluate computational efficiency in Table 8, demonstrating that CaKE achieves better performance while requiring less memory than MEMIT with comparable editing times. This efficiency advantage enables CaKE to scale effectively to larger models, where it maintains superior performance on LLAMA3-70B-Instruct.
@@ 5 Analysis
5.1 Locality Performance
@@ 5.2 Case Analysis
Figure 5 compares the logits of $e_{2}$ and $r_{2}$ at position $t_{2}$ across different editing methods. Here, CaKE generates significantly stronger logits for the bridge entity $e_{2}$ compared to AlphaEdit and MEMIT. This demonstrates CaKE ’s ability to propagate critical information to target positions for subsequent reasoning steps. Similarly, CaKE produces more prominent $r_{2}$ logits, indicating more robust circuit construction and information flow compared to baseline methods.